OpenStudy (anonymous):

Select the correct graph of -x - 2y = 6.

OpenStudy (jim766):

-x - 2y = 6 solve for y +x + x -2y = x + 6 now div by -2 --- -- -- -2 -2 -2 y = -1/2x -3

OpenStudy (jim766):

this line will slope downward down 1 over to the right 2 and will cross the y axis at -3